# Changing the ranking in eigenvector centrality of a weighted graph by small perturbations

### v1

@@ 5.6. Undirected graphs
Undirected graphs lead to some significant simplifications. In this case we have that $A_{\varepsilon}^{\#}=A_{\varepsilon}^{\dagger}$, the Moore-Penrose pseudoinverse, which is symmetric. We define for $A\in\mathbb{R}^{n\times n}$,
$\mathrm{Sym}(A)\big{|}_{ij}:=\frac{a_{ij}+a_{ji}}{2}$
By the well-known property $\mathrm{v}^{\top}A_{\varepsilon}^{\dagger}=0$, and including the symmetry constraint (ic), the gradient (14) simplifies to
@@
(27) $\begin{pmatrix}A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top}&\mathrm{v}\\ \mathrm{v}^{\top}&0\end{pmatrix}=\begin{pmatrix}A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top}&0\\ \mathrm{v}^{\top}&-s\end{pmatrix}\begin{pmatrix}I&(A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top})^{-1}\mathrm{v}\\ 0&1\end{pmatrix}\,,$
@@ 5.7. An illustrative example of an undirected graph
In order to associate to the graph a scaled adiacency matrix $A$ with unit Frobenius norm, we multiply the adiacency matrix whose weights are indicated in Figure 1 by the normalizing factor
@@
The ranking of nodes $\mathrm{v}_{1}$ and $\mathrm{v}_{4}$ can be interchanged by small changes in the weights. This is illustrated in Figure 2. Choosing a suitable small perturbation (of norm $0.0280$) of the normalized adiacency matrix $A$, we obtain the modified Perron eigenvector $\tilde{\mathrm{v}}$
@@
$G_{0}=\left(\begin{array}[]{rrrrrrrrr}0&\!0.336&\!0&\!-0.113&\!0&\!0&\!0&\!0&\!0.645\\ 0.336&\!0&\!-0.029&\!0&\!0&\!0&\!0&\!0&\!0\\ 0&\!-0.029&\!0&\!-0.412&\!0&\!0&\!0&\!0&\!0\\ -0.113&\!0&\!-0.412&\!0&\!-0.419&\!-0.438&\!0&\!0&\!0\\ 0&\!0&\!0&\!-0.418&\!0&\!0&\!0&\!-0.079&\!0\\ 0&\!0&\!0&\!-0.438&\!0&\!0&\!-0.074&\!0&\!0\\ 0&\!0&\!0&\!0&\!0&\!-0.074&\!0&\!0&\!0.188\\ 0&\!0&\!0&\!0&\!-0.073&\!0&\!0&\!0&\!0.109\\ 0.645&\!0&\!0&\!0&\!0&\!0&\!0.188&\!0.109&\!0\\ \end{array}\normalsize\right).$
@@
which preserves the original unperturbed ranking and is characterized by a gap of $0.0146$, significantly different (and with the oppsite sign) from the one in (28).
@@
As we have mentioned, it is possible to consider, instead of the whole set of edges, a subset $\tilde{\mathcal{E}}\subset\mathcal{E}$ and proceed exactly as discussed by simply replacing the projection $P_{\mathcal{E}}$ with $P_{\tilde{\mathcal{E}}}$. As an example we consider again the simple undirected graph with $9$ vertices of Figure 1 and modify only the edges not incident to the two nodes $v_{1}$ and $v_{4}$. The results so obtained are illustrated in Figure 3. We compute the Perron eigenvector $\mathrm{v}$
@@ 5.8. Numerical integration of the gradient system
(29) $\tilde{E}_{k+1}=E_{k}+h\Bigl{(}{}-G_{\varepsilon}^{c}(E_{k})+\gamma_{k}E_{k}\Bigr{)}\quad\hbox{ with }\quad\gamma_{k}=\Big{\langle}E_{k},G_{\varepsilon}^{c}(E_{k})\Big{\rangle}$
where $G_{\varepsilon}^{c}$ has been computed according to (23).
We are not interested in following the exact solution of (22) but mainly to decrease the functional $F_{\varepsilon}^{c}$. This is equivalent to a gradient descent algorithm with line-search. According to this, for the stepsize selection we require that
(30) $F_{\varepsilon}^{c}(\tilde{E}_{k+1})<F_{\varepsilon}^{c}(E_{k}).$
Moreover, we apply an Armijo-like rule. For Equation (22) we have
$\frac{d}{dt}F^{c}_{\varepsilon}(E(t))\Big{|}_{t=t_{k}}=-\varepsilon\Bigl{(}\|G_{\varepsilon}^{c}(E_{k})\|^{2}-\langle G_{\varepsilon}^{c}(E_{k}),E_{k}\rangle^{2}\Bigr{)}:=-g_{k}\leq 0.$
@@
(31) $F_{\varepsilon}^{c}(\tilde{E}_{k+1})\geq F_{\varepsilon}^{c}(E_{k})-(h/\theta)g_{k},$
@@ 5.9. Outer iteration
Let $E(\varepsilon)$ denote the minimizer of the functional $F_{\varepsilon}^{c}$. Generically we expect that for a given perturbation size $\varepsilon<\widetilde{\varepsilon}_{m}$, $E(\varepsilon)$ is smooth, which means that the spectral radius of $A+\varepsilon E$ is a simple eigenvalue. If so, then $f_{c}(\varepsilon):=F_{\varepsilon}^{c}(E(\varepsilon))$ is a smooth function of $\varepsilon$ and we can exploit its regularity to obtain a fast iterative method to converge to $\widetilde{\varepsilon}_{m}$ from the left. Otherwise we can use a bisection technique to approach $\widetilde{\varepsilon}_{m}$.
An example is provided in Figure 4. Since the penalty term is never active the value of $c$ is not relevant here.
The following result provides an inexpensive formula for the computation of the derivative of $f_{c}(\varepsilon)=F_{\varepsilon}^{c}(E(\varepsilon))$, which will be useful in the construction of the outer iteration of the method.
@@
Under Assumption 5.1, the function $f_{c}(\varepsilon)=F_{\varepsilon}^{c}(E(\varepsilon))$ is differentiable and its derivative equals (with ${\phantom{a}{}^{\prime}}=d/d\varepsilon$)
(32) $f_{c}^{\prime}(\varepsilon)=-\|G_{\varepsilon}^{c}(E(\varepsilon))\|$
@@
Differentiating $f_{c}(\varepsilon)=F_{\varepsilon}^{c}\bigl{(}E(\varepsilon)\bigr{)}$ with respect to $\varepsilon$ we obtain
(33) $\hskip-25.60747ptf_{c}^{\prime}(\varepsilon)=\Bigl{\langle}G_{\varepsilon}^{c}(E(\varepsilon)),E(\varepsilon)+\varepsilon E^{\prime}(\varepsilon)\Bigr{\rangle}=\Bigl{\langle}G_{\varepsilon}^{c}(E(\varepsilon)),E(\varepsilon)\Bigr{\rangle}.$
By means of the property no. 3. of minimizers, stated in Theorem 5.4, we can conclude that $E(\varepsilon)\propto G_{\varepsilon}^{c}(E(\varepsilon))$, which yields formula (32). ∎
@@
(34) $\varepsilon_{1}=\varepsilon_{o}-\frac{f_{c}(\varepsilon_{o})}{f_{c}^{\prime}(\varepsilon_{o})},$
@@
Figure 4. Behavior of $f_{c}(\varepsilon)$ for Example in Section 5.7.
@@ 6. Numerical tests
In this section we consider a few numerical experiments illustrating the behavior of our algorithm on a few undirected and directed sparse weighted graphs arising in applications.
@@
Example $2$, Gaertner matrix
This matrix has been contributed to the TAMU Sparse Matrix Collection by Dr. Klaus Gaertner, Weierstrass Institute, Berlin [^1] See https://sparse.tamu.edu/Gaertner. The matrix, described as the adjacency matrix of a weighted undirected graph, is symmetric. However, some of the weights are negative and there are nonzero entries on the main diagonal. To make it non-negative we consider its absolute value. We also set the diagonal to zero. The corresponding graph consists of $n=10774$ nodes. Its pattern is shown in Figure 5.
@@
Finally we set $m=5$. The norm of the perturbation which determines the coalescence of the $5$ leading entries of the Perron eigenvector is $\widetilde{\varepsilon}_{5}=0.0161$, which is about $0.00016\%$. The new leading entries follow:
@@
7. Global stability of the Perron eigenvector
Here we briefly address problem (II), where we replace the constraint (3) with
$\widetilde{a}_{ij}\geq 0\qquad\mbox{for}\quad(i,j)\in\mathcal{E}.$
According to Theorem 3.1, for perturbations $\widehat{A}$ of $A$ such that the perturbed matrix $\widehat{A}$ remains irreducible, there exists an eigenvalue $\lambda_{1}$ of $\widehat{A}$ which is equal to the spectral radius $\rho(\widehat{A})$. In general, however, it is possible that the second largest eigenvalue $\lambda_{2}$ of $\widehat{A}$ coalesces with $\lambda_{1}$, which would imply that $\widetilde{A}$ becomes reducible. The opposite is not true in general, since $\widetilde{A}$ might become reducible with $\lambda_{1}$ remaining simple.
In the case of coalescence of $\lambda_{1}$ and $\lambda_{2}$, the Perron eigenvector would be no longer stably identified, so that if this occurs due to a small perturbation we consider eigenvector centrality to be poorly stable.
Again, let $\widehat{\mathcal{G}}=(\mathcal{V},\mathcal{E},\widehat{A})$ denote a graph with the same vertices and the same edges of $\mathcal{G}$. We propose - both for undirected and directed graphs - the following stability measure for eigenvector centrality, which we call stability radius:
$\omega(\mathcal{G})=$ $\>\inf_{\widehat{\mathcal{G}}}{\rm dist}\left(\mathcal{G},\widehat{\mathcal{G}}\right)=\inf_{\widehat{A}}\|A-\widehat{A}\|$
under the condition that the Perron eigenvalue of $\widehat{A}$ is not simple and under the constraints $\widehat{a}_{ij}\geq 0$ for $(i,j)\in\mathcal{E}$ and $\widehat{a}_{ij}=0$ for $(i,j)\notin\mathcal{E}$.
The choice of the norm is naturally important; here we use the Frobenius norm for a mainly practical reason: being an inner product norm, we are able to compute the gradient of a suitable functional associated to the problem and proceed in a similar fashion as the one considered in previous section, when we addressed the problem of Section 3.1.
Note that for a strongly connected graph we have the natural lower bound
$\omega(\mathcal{G})<\min\limits_{(i,j)\in\mathcal{E}}a_{ij}.$
In the case of an undirected graph, by the Hoffman–Wielandt theorem [24, p. 368], it is immediate to get the lower bound which relates the stability radius with the spectral gap of $A$ (see [1]).
Theorem 7.1.
Assume $\mathcal{G}$ is an undirected graph to which is associated the weighted adjacency matrix $A$. Then, denoting by $\lambda_{1}\geq\lambda_{2}$, the two largest eigenvalues of $A$, it holds
$\omega(\mathcal{G})\geq(\lambda_{2}-\lambda_{1})/\sqrt{2}.$
Proof.
The proof is based on the observation that if we consider the canonical decomposition of $A=U\Lambda U^{\top}$ and replace $\Lambda$ by $\widehat{\Lambda}$ which differs from $\Lambda$ only in the first two eigenvalues which are set equal to $(\lambda_{1}+\lambda_{2})/2$, one gets a symmetric matrix $\widehat{A}$ with coalescent largest two eigenvalues at a distance $(\lambda_{2}-\lambda_{1})/\sqrt{2}$ from $A$.
Then, considering an arbitrary symmetric matrix $\widetilde{A}$ with coalescent two largest eigenvalues it is straightforward to prove that its distance from $A$ is larger that $(\lambda_{2}-\lambda_{1})/\sqrt{2}$. ∎
If the perturbed matrix $\widehat{A}$ becomes reducible but the Perron eigenvalue remains simple, one may either stop the procedure or continue separately with the two subgraphs.
The stability of the Perron eigenvector is determined by the smallest perturbation of the weighted adjacency matrix $A$ which causes the coaescence of the largest eigenvalue $\lambda_{1}$ with another eigenvalue, This may occur through the coalescence of $\lambda_{1}$ either with a real eigenvalue or with a complex conjugate pair with vanishing imaginary part. We mention that in the special case of irreducible (row) stochastic matrices, the stability of the left Perron vector (stationary probability distribution) under perturbations has been investigated by various authors; see, e.g., [8, 9, 12, 26] and references therein. In these papers the sensitivity of the stationary distribution under perturbations that preserve the irreducibility and stochasticity of the transition probability matrix is expressed in terms of various quantities including the group inverse, the mean passage times, and the so-called ergodicity coefficient.
Also in this case we follow the two-level method considered previously in Section 3.2, to iteratively minimize the functional (for given $\varepsilon$)
(35) $S_{\varepsilon}(E)=\frac{1}{2}\Bigl{(}\lambda_{1}\left(A+\varepsilon E\right)-{\mbox{\rm Re}}\lambda_{2}\left(A+\varepsilon E\right)\Bigr{)}^{2},$
being $\lambda_{2}$ the closest eigenvalue to $\lambda_{1}$ (conventionally with positive imaginary part), under the constraints of unit Frobenius norm of $E$, of non-negativity $A+\varepsilon E\geq 0$ and the preservation of the sparsity pattern of $A$.
In the outer iteration we compute the optimal $\varepsilon$, denoted $\varepsilon_{\star}$,
(36) $\varepsilon_{\star}\longrightarrow\arg\min\limits_{\varepsilon>0}S_{\varepsilon}\left(E(\varepsilon)\right)=0$
by a combined Newton-bisection method. The existence of the minimizer is obtained by compactness arguments. We then have
$\omega(\mathcal{G})=\varepsilon_{\star}.$
Making use of Lemma 4.1, we have the following result:
Lemma 7.1 (Norm unconstrained gradient of the functional).
Assume that the non-negative matrix $A+\varepsilon E$ is irreducible; then, in the space of real matrices with sparsity pattern ${\mathcal{E}}$ we have
$\frac{d}{dt}S_{\varepsilon}\left(E(t)\right)=\varepsilon\Big{\langle}H_{\varepsilon}(E),\dot{E}\Big{\rangle},$
where $H_{\varepsilon}(E)$ has the sparsity pattern determined by the set of edges $\mathcal{E}$ and is given by
(37) $H_{\varepsilon}(E)=P_{\mathcal{E}}\Bigl{(}\kappa_{1}x_{1}y_{1}^{\top}-\kappa_{2}{\mbox{\rm Re}}\left(x_{2}y_{2}^{*}\right)\Bigr{)}$
where $x_{1}$ and $y_{1}$ are the left and right eigenvectors of $A+\varepsilon E$ associated to the Perron eigenvalue $\lambda_{1}$, normalized according to (8), and $x_{2}$ and $y_{2}$ are the left and right eigenvectors associated to $\lambda_{2}$, and $\kappa_{1}=1/(x_{1}^{\top}y_{1}),\kappa_{2}=1/(x_{2}^{*}y_{2})$.
Proof.
The proof makes use of Lemma 4.1 and follows similar steps to the proof of Lemma 5.1. ∎
Note that in the undirected case, i.e. when $A$ and $E$ are symmetric, $\kappa_{1}=\kappa_{2}=1$ and $x_{2}=y_{2}$ is a real vector, so that
$H_{\varepsilon}(E)=P_{\mathcal{E}}\Bigl{(}x_{1}x_{1}^{\top}-x_{2}x_{2}^{\top}\Bigr{)}$
The constrained gradient system
The positivity constraint for $A+\varepsilon E$ can be treated again by penalization, as described in Section 5.5; hence we get the structure and norm constrained gradient system for the penalized functional $S_{\varepsilon}\left(E(t)\right)+\displaystyle\frac{c}{\varepsilon}P_{\varepsilon}\left(E(t)\right)$,
(38) $\dot{E}=-\left(H_{\varepsilon}(E)+\frac{c}{\varepsilon}\,\nabla P_{\varepsilon}(E)\right)+\eta E\quad\hbox{ with }\quad\gamma=\langle E,H_{\varepsilon}(E)+\frac{c}{\varepsilon}\,\nabla P_{\varepsilon}(E)\rangle.$
with $H_{\varepsilon}(E)$ given by (37).
Note that as $\lambda_{2}\to\lambda_{1}$ the terms $\kappa_{1}$ and $\kappa_{2}$ diverge to +$\infty$+ ∞, which implies that we have to use a suitable tolerance to establish the occurrence of coalescence.
7.1. The outer iteration
The outer iteration is analogous to the one considered previously in Section 5.9; in particular, after considering the same penalty function (21), Lemma 5.2 applies verbatim.

### v2

@@ 6. Undirected graphs
Undirected graphs lead to some significant simplifications. In this case we have that $A_{\varepsilon}^{\#}=A_{\varepsilon}^{\dagger}$, the Moore-Penrose pseudoinverse, which is symmetric.
Recalling (18), we have
$G_{\varepsilon}(E)$ $=$ ${}-P_{\mathcal{E}}\left(\sum_{k=1}^{m}\left(\mathrm{v}_{i_{k}}-\langle\mathrm{v}\rangle_{m}\right)\mathrm{Sym}\Bigl{(}A_{\varepsilon}^{\dagger}\left({\rm u}_{i_{k}}-\bar{\rm u}\right)\mathrm{v}^{\top}\Bigr{)}\right)$
@@
${\small\begin{pmatrix}A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top}&\mathrm{v}\\ \mathrm{v}^{\top}&0\end{pmatrix}=\begin{pmatrix}A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top}&0\\ \mathrm{v}^{\top}&-s\end{pmatrix}\begin{pmatrix}I&(A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top})^{-1}\mathrm{v}\\ 0&1\end{pmatrix}}$
@@ 6.1. An illustrative example of an undirected graph
In order to associate to the graph a scaled adjacency matrix $A$ with unit Frobenius norm, we multiply the adjacency matrix whose weights are indicated in Figure 1 by the normalizing factor
@@
The ranking of nodes $\mathrm{v}_{1}$ and $\mathrm{v}_{4}$ can be interchanged by small changes in the weights. This is illustrated in Figure 2. Choosing a suitable small perturbation (of norm $0.0280$) of the normalized adjacency matrix $A$, we obtain the modified Perron eigenvector $\tilde{\mathrm{v}}$
@@
$G_{0}=\left(\begin{array}[]{rrrrrrrrr}0&\!0.336&\!0&\!-0.113&\!0&\!0&\!0&\!0&\!0.645\\ 0.336&\!0&\!-0.029&\!0&\!0&\!0&\!0&\!0&\!0\\ 0&\!-0.029&\!0&\!-0.412&\!0&\!0&\!0&\!0&\!0\\ -0.113&\!0&\!-0.412&\!0&\!-0.418&\!-0.438&\!0&\!0&\!0\\ 0&\!0&\!0&\!-0.418&\!0&\!0&\!0&\!-0.073&\!0\\ 0&\!0&\!0&\!-0.438&\!0&\!0&\!-0.074&\!0&\!0\\ 0&\!0&\!0&\!0&\!0&\!-0.074&\!0&\!0&\!0.188\\ 0&\!0&\!0&\!0&\!-0.073&\!0&\!0&\!0&\!0.109\\ 0.645&\!0&\!0&\!0&\!0&\!0&\!0.188&\!0.109&\!0\\ \end{array}\normalsize\right).$
@@
which preserves the original unperturbed ranking and is characterized by a gap of $0.0146$, significantly different (and with the opposite sign) from the one in (33).
@@
As already mentioned in the discussion following Definition 5.1, it is possible to consider, instead of the whole set of edges, a subset ${\mathcal{E}}^{\prime}\subset\mathcal{E}$ and proceed exactly as discussed by simply replacing the projection $P_{\mathcal{E}}$ with $P_{\mathcal{E}^{\prime}}$. As an example we consider again the simple undirected graph with $9$ vertices of Figure 1 and modify only the edges not incident to the two nodes $v_{1}$ and $v_{4}$. The results so obtained are illustrated in Figure 3. We compute the Perron eigenvector $\mathrm{v}$
@@ 6.2. Numerical integration of the gradient system
(34) $\tilde{E}_{k+1}=E_{k}+hP^{+}\Bigl{(}{}-G_{\varepsilon}(E_{k})+\gamma_{k}E_{k}\Bigr{)}\quad\hbox{ with }\quad\gamma_{k}=\Big{\langle}P^{+}E_{k},P^{+}G_{\varepsilon}(E_{k})\Big{\rangle}$
where $G_{\varepsilon}$ has been computed according to (16).
We are not interested in following the exact solution of (25) but mainly to decrease the functional $F_{\varepsilon}$. This is equivalent to a gradient descent algorithm with line-search. According to this, for the stepsize selection we require that
(35) $F_{\varepsilon}(\tilde{E}_{k+1})<F_{\varepsilon}(E_{k}).$
Moreover, we apply an Armijo-like rule. For Equation (25) we have (recalling that $P^{+}$ is just the identity unless inequality constraints activate)
$\frac{d}{dt}F_{\varepsilon}(E(t))\Big{|}_{t=t_{k}}=-\varepsilon\Bigl{(}\|P^{+}G_{\varepsilon}(E_{k})\|^{2}-\langle P^{+}G_{\varepsilon}(E_{k}),P^{+}E_{k}\rangle^{2}\Bigr{)}:=-g_{k}\leq 0.$
@@
(36) $F_{\varepsilon}(\tilde{E}_{k+1})\geq F_{\varepsilon}(E_{k})-(h/\theta)g_{k},$
@@
In a numerical solution of the gradient system (for example using the Euler-based scheme (34)), we have to monitor the sets of edges where $a_{ij}+\varepsilon e_{ij}=\delta$ and among those we must further track those edges where the sign of $-g_{ij}+\gamma e_{ij}$ changes (where $g_{ij}$ is the $(i,j)$-th entry of the matrix $G_{\varepsilon}(E)$ and $e_{ij}$ is the $(i,j)$-th entry of the matrix $E$). When the active set $\{(i,j)\,:\,a_{ij}+\varepsilon e_{ij}=\delta\}$ is changed, then $\gamma$ also changes.
@@ 6.3. Outer iteration
Due to the fact that in our experiments inequality constraints do not activate, we restrict our analysis of the outer iteration to this setting (where $P^{+}$ is the identical operator). A more elaborated explicit formula can be also deduced in the general case.
The case of inactive inequality constraints
Let $E(\varepsilon)$ denote the minimizer of the functional $F_{\varepsilon}$. Generically we expect that for a given perturbation size $\varepsilon<\widetilde{\varepsilon}_{m}$, $E(\varepsilon)$ is smooth, which means that the spectral radius of $A+\varepsilon E$ is a simple eigenvalue. If so, then $f(\varepsilon):=F_{\varepsilon}(E(\varepsilon))$ is a smooth function of $\varepsilon$ and we can exploit its regularity to obtain a fast iterative method to converge to $\widetilde{\varepsilon}_{m}$ from the left. Otherwise we can use a bisection technique to approach $\widetilde{\varepsilon}_{m}$.
An example is provided in Figure 4.
The following result provides an inexpensive formula for the computation of the derivative of $f(\varepsilon)=F_{\varepsilon}(E(\varepsilon))$, which will be useful in the construction of the outer iteration of the method.
@@
Under Assumption 6.1, the function $f(\varepsilon)=F_{\varepsilon}(E(\varepsilon))$ is differentiable and its derivative equals (with ${\phantom{a}{}^{\prime}}=d/d\varepsilon$)
(37) $f^{\prime}(\varepsilon)=-\|G_{\varepsilon}(E(\varepsilon))\|$
@@
Differentiating $f(\varepsilon)=F_{\varepsilon}\bigl{(}E(\varepsilon)\bigr{)}$ with respect to $\varepsilon$ we obtain
(38) $\hskip-25.60747ptf^{\prime}(\varepsilon)=\Bigl{\langle}G_{\varepsilon}(E(\varepsilon)),E(\varepsilon)+\varepsilon E^{\prime}(\varepsilon)\Bigr{\rangle}=\Bigl{\langle}G_{\varepsilon}(E(\varepsilon)),E(\varepsilon)\Bigr{\rangle}.$
By means of the property no. 3. of minimizers, stated in Theorem 5.1, we can conclude that $E(\varepsilon)\propto G_{\varepsilon}(E(\varepsilon))$, which yields formula (37). ∎
Newton method
@@
(39) $\varepsilon_{1}=\varepsilon_{o}-\frac{f(\varepsilon_{o})}{f^{\prime}(\varepsilon_{o})},$
@@
Figure 4. Behavior of $f(\varepsilon)$ for Example in Section 6.1.
@@ 7. Numerical tests
In this section we present numerical experiments illustrating the behavior of our algorithm on a few undirected and directed sparse weighted graphs arising in applications. We also perform a comparison with more standard approaches for solving constrained optimization problems.
@@ 7.0.1. Example 1, matrix $A_{27}$
Comparison with Matlab routines
In a first experiment we have applied the Matlab routine fmincon to the problem formulation (4)–(5) for the case $m=2$. Here we use fmincon as a black box, without providing the analytic form of the gradient (the derivation of which is among the results of this paper) to the routine. Using standard default values we have not been able to obtain a correct solution, hence it has been necessary to increase the number of function evaluations allowed. Table 2 shows the results of the routine with an interior point algorithm, which is the one that performs best. The resulting solution is very close to the one we found, with coalescent entries #$\#111$ and #$\#315$,
$\mathrm{v}_{111}=0.{\bf 33}2\ldots,\qquad\mathrm{v}_{315}=0.{\bf 33}2\ldots.$
In Table 2, $k$ stands for the iteration number, $x$ is the vector containing the nonzero entries of the upper triangular part of the weighted adiacency matrix $A$, $g(x)$ is the constraint, $\mathrm{v}_{1}-\mathrm{v}_{2}$ being $\mathrm{v}(x)$ the Perron eigenvector of the matrix $B=A-X$, where $X$ is the perturbation associated to $x$. Finally, $f(x)=\sqrt{2}\|x\|=\|X\|_{F}$ and #$\#fe$ indicates the number of function evaluations. The overall CPU time is about $80$ minutes.
$k$#\## fe $f(x_{k})$ $g(x_{k})$
 0 1.4750e+03 6.9396e-04 8.4470e-02
1 2.9500e+03 2.4405e+01 7.6330e-02
6 1.0325e+04 3.8498e+01 4.3180e-04
11 1.7700e+04 7.3944e+00 2.0100e-03
21 3.2450e+04 6.7381e-01 2.5690e-03
31 4.7206e+04 6.6791e-01 1.7290e-06
41 6.1970e+04 5.1159e-02 2.3090e-02
51 7.6743e+04 4.9650e-02 2.9350e-04
61 9.1510e+04 4.0801e-02 1.0950e-02
71 1.0629e+05 1.4522e-02 4.7660e-03
81 1.2109e+05 9.5850e-03 3.9870e-05
91 1.3589e+05 9.2726e-03 3.5810e-10
96 1.4329e+05 9.2726e-03 5.8570e-10
Table 2. Reported behavior of the Matlab algorithm.
Observing Table 2, we see that in the beginning $x$ gets quite large with respect to its final value ($3.8\cdot 10^{1}$ versus $9.27\cdot 10^{-3}$). This very likely implies activation of inequality constraints and makes the convergence process slower. For this reason, trust region techniques would likely improve the process. The initial value $x_{0}$ provided to the algorithm is quite small (as $f(x_{0})$ indicates), but the algorithm makes a large jump to a vector $x_{1}$ of much higher norm, which is due to the violation of the constraint. Starting from a consistent value would help, but requires additional computation.
Next we describe the results of a second experiment, with analytic gradients provided to the algorithm. Following the notation in Section 3.2, we have provided to the algorithm, the exact formula for the gradient of the constraint $u_{i_{1}}-u_{i_{2}}=0$, which is obtained by the same arguments used in Lemma 5.1 and is proportional to
$P_{\mathcal{E}}(\mathrm{Sym}(B^{\dagger}(u_{i_{1}}-u_{i_{2}})\mathrm{v}^{\top}).$
The norm of the perturbation which determines the coalescence of the two main entries of the Perron eigenvector is $\widetilde{\varepsilon}_{2}=0.01023234$, which is about $1\%$. The two leading entries after the perturbation are
$\mathrm{v}_{315}\approx\mathrm{v}_{636}=0.20245843.$
This indicates a different local optimum with respect to the previous experiment (which was the same as the one we found). However, they are very close and both indicate a strong sensitivity of the ranking for this graph. Table 3 shows the results of the run.
$k$#\## fe $f(x_{k})$ $g(x_{k})$
 0 1.475e+03 1.002577e-03 8.433e-02
1 2.950e+03 2.329900e+01 5.114e-02
21 3.2464e+04 4.892883e+00 5.230e-03
41 6.1970e+04 2.175929e-01 5.546e-03
61 9.1493e+04 1.814369e-01 6.202e-06
81 1.21020e+05 1.786729e-02 1.009e-03
101 1.50620e+05 1.025323e-02 9.777e-08
121 1.80195e+05 1.023304e-02 3.949e-07
141 2.09771e+05 1.023236e-02 1.121e-07
147 2.18661e+05 1.023233e-02 2.660e-09
Table 3. Reported behavior of the Matlab algorithm provided by the analytic gradient of the constraint.
The overall CPU time is about $351$ seconds, which shows a significant improvement with respect to the previous experiment, although still far from the CPU time required by our approach.
7.0.2. Example $2$, matrix nopoly
This matrix has been contributed to the TAMU Sparse Matrix Collection by Dr. Klaus Gaertner, Weierstrass Institute, Berlin [^1] See https://sparse.tamu.edu/Gaertner. The matrix, described as the adjacency matrix of a weighted undirected graph, is symmetric. However, some of the weights are negative. To make it non-negative we consider its absolute value. The corresponding graph consists of $n=10774$ nodes. Its pattern is shown in Figure 5. We do not perform a comparison with Matlab’s optimization tools since the size of the matrix makes this prohibitively expensive.
@@ 7.0.2. Example $2$, matrix nopoly
Next we set $m=5$. The norm of the perturbation which determines the coalescence of the $5$ leading entries of the Perron eigenvector is $\widetilde{\varepsilon}_{5}=0.0161$, which is about $0.00016\%$. The new leading entries follow: